What is the volume of a gas if the pressure is 3 atm, the number of moles is 3, the temperature is 300 K and the Universal Gas Constant is 0.0821


1
Expert's answer
2021-05-10T01:25:54-0400

V= nRT/P = 3×0.0821×300/3 = 24.6 L

Hence, the approximate volume occupied will be 24.6 L.